# Style transfer between Microscopy and Magnetic Resonance Imaging via Generative Adversarial Network in small sample size settings

### v2

@@ 2.1 Dataset
We use a multimodal dataset previously acquired for which ethical approval was previously obtained by the authors [8]. More specifically, in this dataset MRI in post-mortem tissue has been performed and integrated with microscopy. Post-mortem scans provided whole human brain coverage with voxel sizes of 100-500 µm. For this paper, the MRIs and glial fibrillary acidic protein maps (GFAP) from the ‘Digital Anatomist’ have been used. More specifically, the first volume of the diffusion MRI (400 µm isotropic at 7T), alongside parallelly coregistered histology (0.25 µm in-plane) in three human corpus callosum specimens were used. For more information on the data acquisition are available [9]. The choice of using the volume was given by the fact that no traditional T1 or T2 sequence was available. It can be hypothesized that with such unavailable modalities, results should be superior. We utilized 3 samples of GFAPs and 3 samples of MRI scans of corpus callosum. The dataset is small and contains 5 training examples of whole slice view paired images and 3 examples of paired images for testing. Additionally, large differences in contrast, resolution, and type of details occur between MRI and histology. Making this a non-trivial case of style-transfer between images.
@@ 2.2 Data preprocessing
Image registration: To align MRI and histology slices to a common coordinate system, intensity-based affine image registration was implemented using built-in MATLAB functions from the ‘Register Multimodal MRI Images’ toolbox. Sagittal slices of volumetric MRI brain representations exported from ITK-SNAP [11] medical image viewer were loaded as moving images and GFAPs as fixed. Downsampled GFAPs were of size around 6000x10000 and MRIs - 98x128. The mean Dice score evaluating the obtained alignment between fixed and moving is 0.93.
@@ 2.3 Network architecture
A previously defined cGAN architecture has been adapted to translate MRIs to GFAPs [13]. Image-to-image translation aims at learning the mapping between the input image and an output image using a set of registered pairs of images. Given the nature of our dataset, the cGAN model was trained on paired data (each histology section paired and registered with MRI slices). The model has two architectures, one for the generator and the other for the discriminator, specifically encoder-decoder U-net and patchGAN. Discriminator’s patchGAN architecture includes a number of transpose convolutional blocks. It examines an NxN section of an image to determine whether it is real or fake [13]. Conditional GANs learn the mapping from an observed image x and a random noise vector z to y, G: . G - generator - is trained to produce outputs that cannot be distinguished from real images by the discriminator. A discriminator, D, is adversarially trained to do as well as possible at spotting the generator’s ”fakes”. Load sizes chosen for 3 training experiments were: 4096, 1024, and 256. The objective adversarial loss can be expressed as:
with the corresponding discriminator of . Then, using jointly the L1 distance and considering the random noise vector,
leads to the GAN objective:
This was implemented as cGAN in Python inspired by the work of Isola et al. [13] based on Pytorch.

### v3

@@ 2.1 Dataset
We use a multimodal dataset previously acquired by Tendler et al. [8], for which ethical approval was obtained by the original authors. For this study, we selected MRIs and glial fibrillary acidic protein (GFAP) maps from the Digital Anatomist resource. Specifically, the first volume of the diffusion MRI (400 µm isotropic at 7 T) and the corresponding registered histology (0.25 µm in-plane) from three human corpus callosum specimens were used. For more information on the data acquisition are available [8, 9]. The choice of using the volume was given by the fact that no traditional T1 or T2 sequence was available. It can be hypothesized that with such unavailable modalities, results should be superior. The dataset is small and contains 5 training examples of whole slice view paired images and 3 examples of paired images for testing. Additionally, large differences in contrast, resolution, and type of details occur between MRI and histology. Making this a non-trivial case of style-transfer between images.
@@ 2.2 Data preprocessing
Image registration: To align MRI and histology slices to a common coordinate system, intensity-based affine image registration was implemented using built-in MATLAB functions from the ‘Register Multimodal MRI Images’ toolbox. Sagittal slices of volumetric MRI brain representations exported from ITK-SNAP [11] medical image viewer were loaded as moving images and GFAP maps as fixed. Downsampled GFAP maps were of size around 6000x10000 and MRIs - 98x128. The mean Dice score evaluating the obtained alignment between fixed and moving is 0.93.
@@ 2.3 Network architecture
Image-to-image translation aims at learning the mapping between the input image and an output image using a set of registered pairs of images. Given the nature of our dataset, a cGAN model was trained on paired data (each histology section paired and registered with MRI slices). We adapted the publicly available PyTorch implementation of pix2pix, a cGAN model, by Isola et al. [13] to translate MRIs to histology, incorporating an image registration preprocessing step and hyperparameter tuning. The model has two architectures, one for the generator and the other for the discriminator, specifically encoder-decoder U-net and patchGAN. Discriminator’s patchGAN architecture includes a number of transpose convolutional blocks. It examines an NxN section of an image to determine whether it is real or fake [13]. Conditional GANs learn the mapping from an observed image x and a random noise vector z to y, G: . G - generator - is trained to produce outputs that cannot be distinguished from real images by the discriminator. A discriminator, D, is adversarially trained to do as well as possible at spotting the generator’s ”fakes”. Load sizes chosen for 3 training experiments were: 4096, 1024, and 256. The adversarial loss for the conditional GAN is defined as:
where denotes the discriminator corresponding to the generator . Then, using jointly the L1 distance and considering the random noise vector,
leads to the final objective: